# Asymmetric Effects Underlying Dynamic Heterogeneity in Miscible Blends of Poly(methyl methacrylate) with Poly(ethylene oxide)

### v1

@@ Abstract
The emergence of spatially variable local dynamics, or dynamic heterogeneity, is common in multicomponent polymer systems. Such heterogeneity is understood to arise from differences between the intrinsic dynamical fluctuations associated with one component versus another. However, the nature of the dynamic coupling between these components and how it depends on composition, temperature, and environmental fluctuations is not fully understood. Here, we use molecular dynamics simulations to characterize nanoscale dynamic heterogeneity in miscible blends of polyethylene oxide (PEO) and polymethyl methacrylate (PMMA) as a function of both temperature and blend composition. Probed over timescales of 100 ps, local PMMA segmental dynamics in blends align with neat PMMA when normalized by $T_{\text{g}}$, whereas PEO exhibits enhanced mobility caused by free-volume effects. The effects of dynamical coupling are found to be asymmetric between the extent of enhancement to PMMA and suppression of PEO dynamics based on analysis of local environment composition within blends. Asymmetric effects in the melt state are also identified over longer timescales according to a Rouse mode analysis over larger sub-chains for each species. These results provide fundamental insights into how dynamic heterogeneity manifests at the nanoscale, across conditions and compositions in miscible polymer blends. They establish a foundation for exploring whether such asymmetries are generalizable and how dynamic heterogeneity can be tuned through temperature, composition, and morphology.
@@
[…]
@@ 1 Introduction
Dynamic heterogeneity is known to be significant in blends of polyethylene oxide (PEO) and polymethyl methacrylate (PMMA). While these polymers are miscible,22, 23 their glass transition temperatures ($T_{\text{g}}$) differ by approximately 180 K.24, 25 This dissimilarity gives rise to pronounced differences in segmental dynamics when the polymers are blended, with relaxation times differing by up to 12 orders of magnitude.24, 26 Such contrast in mobilities has implications for applications like solid polymer electrolytes,27, 28, 29, 30, 31 where ionic conductivity is sensitive to local segmental fluctuations.32, 33, 34 More broadly, PEO/PMMA blends serve as model systems for investigating dynamic coupling between components at the nanoscale, with signatures of heterogeneity accessible in both simulations and experiments.26, 35, 36, 37
@@
Numerous experimental studies have examined the dynamics of PEO and PMMA in blends under varying conditions. Quasi-elastic neutron scattering (QENS) on blends containing up to 30 wtPEO has shown that the segmental mobility of PMMA, on length scales up to 11 Å, is primarily governed by the temperature difference between the system and the glass transition temperature of the blend, (i.e., $T-T_{\text{g}}$).38, 39, 40 In contrast, PEO dynamics are strongly influenced by interactions with PMMA, exhibiting distinct behavior at short (<1 nm) and longer (>1nm) length scales.41, 35 At short length scales, QENS,42, 36 nuclear magnetic resonance (NMR),24, 26 and neutron spin echo (NSE) spectroscopy43 reveal narrowing distributions of segmental relaxation times across a range of compositions and temperatures. These effects are attributed to self-concentration phenomena42 and local confinement by the stiffer PMMA matrix, with a dependence on local free volume.24, 26, 36
@@
On longer length scales, PEO dynamics have been investigated using infrared dichroism and birefringence,44 QENS,41, 36 and NSE.35 These measurements consistently indicate a pronounced slowdown of PEO segmental motion in blends with low PEO content.44, 41, 35, 36 Unlike the behavior observed at shorter scales, the dynamics at these larger scales are characterized by a broad distribution of relaxation times,36 which is largely attributed to long-range concentration fluctuations.44 Collectively, these observations highlight the presence of dynamic heterogeneity across multiple length scales in PEO/PMMA blends and suggest several underlying mechanisms; however, direct inference of the molecular-level phenomena remains elusive through experimental characterization alone.
@@
To complement experimental observations, several theoretical models have been developed to explain dynamic heterogeneity in PEO/PMMA blends. The Lodge-McLeish (LM) model, also referred to as the chain connectivity model, quantifies the influence of local self-concentration on polymer dynamics. The LM model predicts the segmental dynamics of each polymer component using the concentration within a cooperative volume centered on a monomer. The relevant length scale used to determine the size of the cooperative volume is the Kuhn length of each polymer component. It successfully predicts PEO relaxation times in blends containing 10-30 wtPEO taken from QENS measurements for large spatial scales ($q$=0.69 Å-1, approximately 18 Å) but breaks down for smaller spatial scales ($q$=1.3 Å-1, approximately 10 Å).36 When the self-concentration is allowed to vary as a fitting parameter, the LM model reasonably fits PEO relaxation times measured by QENS36 and NMR,26 as well as terminal PMMA relaxation times.45 However, allowing the self-concentration to vary obfuscates the theoretical foundations of and insights from the model.
Another model uses mesoscale concentration fluctuations with length scales between 1 nm and 1 $\mu$m via the generalized entropy theory of glass formation, the lattice cluster theory of blend thermodynamics, and the Kirkwood-Buff theory of concentration fluctuations to predict structural relaxation times of dynamically asymmetric miscible polymers in blends.46, 47 It can qualitatively fit the separated relaxation times of PEO and PMMA in blends but cannot account for the chemical specificity required for fitting quantitative behaviors.46, 47
@@
Finally, a coupling model48, 49, 50, 51, 52 and the generalized Langevin equation framework53, 54, 55 have been developed to describe a crossover time that separates single-chain PEO dynamics at short length scales from many-chain coupled PEO dynamics at long length scales. This framework has been used to explain unexpected phenomena and properties that arise specifically in dynamically asymmetric miscible polymer blends, such as different $T_{\text{g}}$ values for each component and the breakdown of the time-temperature superposition.50 Additionally, the coupling model closely captures PEO segmental dynamics for experimental QENS measurements using a momentum transfer $q$ value between 1 and 2 Å-1 in blends with 10-30 wtPEO.36 These theories qualitatively capture and rationalize trends in blended PEO dynamics up to mesoscale length scales. However, like with experimental observations, they do not conclusively illustrate the molecular-level phenomena that contribute to dynamic heterogeneity.
@@
Molecular dynamics (MD) simulations have helped to elucidate the microscopic origins of dynamic heterogeneity in polymer blends under certain conditions. Analysis of all-atom simulations with blend compositions of 10-30 wtPEO suggest the presence of multiple populations of PEO dynamics in blends which may be caused by confinement effects of PEO in a rigid matrix of PMMA. This is deduced from van Hove self-correlation functions of hydrogen atoms from PEO, which show a double-peak structure at 400 K with an unmoving second peak41 and PEO relaxation times that are broadly distributed.56 Monitoring mean-square deviations of atoms in PEO and PMMA also suggests highly disparate local cage sizes, as the magnitude of distance traveled in the ballistic regime is much higher for PEO than PMMA.41, 56 Additionally, PEO segmental relaxation times have been found to be more stretched upon blending than those of PMMA.56 Rouse analyses based on simulations of 20 wtPEO have illustrated the non-Gaussianity of the distribution of PEO atomic displacements in blends57, suggesting that local PEO segmental motion for wavelengths on the order of the size of a monomer is not strongly affected by blending but is more strongly affected at larger wavelengths.37 These results provide molecular-level insights into dynamic heterogeneity in PEO/PMMA blends, albeit using a varied set of compositions, temperatures, and characterization methods.
@@
In this work, we use atomistic MD simulations to systematically characterize how dynamic heterogeneity manifests at the nanoscale in PEO/PMMA blends. A central objective is to examine how the strength of dynamic coupling between the two components depends on temperature and blend composition. Although prior studies have explored the dynamic behavior of PEO/PMMA blends, inconsistencies in system specifications (e.g., blend composition, molecular weight, and force field) as well as differences in thermodynamic conditions have made it difficult to identify clear trends. To address this, we comprehensively investigate behavior over a full range of blend compositions and at temperatures above, between, and below the simulated apparent $T_{\text{g}}$ of both polymers. Dynamics are characterized by means of both local segmental fluctuations as well as segmental relaxation timescales. We find that blending induces asymmetric and composition-dependent changes in both local and collective dynamics, governed by temperature, local composition, and local structure. These findings help to clarify what PEO-PMMA interactions influence dynamic heterogeneity and motivate further investigation into their generality and implications for macroscopic properties and functional performance.
@@ 2.1.1 General simulation protocols
All MD simulations are performed using LAMMPS (ver 29, Sep 2021).58 Systems are modeled using the all-atom optimized potentials for liquid simulations (OPLS-AA)59 force field. Real-space non-bonded interactions are truncated at 12 Å. Long-range electrostatics are handled using the particle-particle-particle-mesh Ewald summation method with a $10^{-4}$ convergence accuracy.60, 61 Equations of motion are integrated using a velocity-Verlet integration scheme and a 1 fs timestep. Periodic boundary conditions are used in all three dimensions. Temperature and pressure are controlled using a Nosé-Hoover thermostat and barostat with damping constants of 100 fs and 2000 fs, respectively.
@@ 2.1.2 Polymer chain generation
Each PEO chain consists of 75 monomers, and each PMMA chain consists of 33 monomers. As a result, chains for both polymers possess molecular weights of approximately 33,000 g/mol. Initial chain configurations are generated to match the expected average squared end-to-end distance, $\langle R^{2}\rangle=Ll_{k}$, where $L$ is the contour length and $l_{k}$ is the Kuhn length of the polymer. The contour length $L$ is estimated from the fully extended conformation of each polymer using geometry optimization in Avogadro 1.2.0 with the steepest descent algorithm and the UFF force field;62, 63 this yields values of 326.1 Å for PEO and 102.5 Å for PMMA. The Kuhn lengths used are 8.2 Å for PEO and 13.8 Å for PMMA.64 Chains are constructed by sequentially adding monomers, with each new monomer positioned based on a randomly sampled dihedral angle. For PEO, dihedral angles are drawn from a uniform distribution between $-0.65$ and $0.65$ radians; for PMMA, the range is $-0.4$ to $0.4$ radians. For each system described below, this process is repeated until the required number of independent chain configurations within a threshold of 5 Å of the target $\langle R^{2}\rangle$ is obtained. These configurations are used as described in Section 2.1.3.
@@ 2.1.3 System preparation
For each system, an initial configuration is generated by randomly placing each pre-constructed chain into a 60 Å $\times$ 60 Å $\times$ 60 Å simulation box with random positions and orientations. A brief energy minimization is then performed where the system is simulated in the microcanonical (NVE) ensemble with a constrained maximum distance of 0.005 Å moved per timestep; this minimization takes place for 0.05 ns. After, the constrained maximum distance is increased to 0.1 Å  and the system is simulated for 0.5 ns. This procedure resolves unfavorable atomic overlaps introduced by the random packing procedure. Following minimization, initial velocities for all particles are randomly generated with a uniform distribution at 300 K. The system undergoes 1 ns of simulation in the NVE ensemble, followed by 1 ns in the canonical (NVT) ensemble at 300 K. Subsequently, a barostat is introduced to maintain the pressure at 1 atm, and the system is heated from 300 K to 700 K at a rate of 80 K/ns. After reaching 700 K, the system is equilibrated in the isothermal–isobaric (NPT) ensemble for 50 ns. Finally, the system is cooled from 700 K to 100 K at a rate of 10 K/ns at constant pressure.
The cooling trajectories are used for analysis of glass transition temperatures (see Section 2.2.1). Simulation configurations are also recorded specifically around the temperatures of 500 K, 360 K, and 220 K. From these configurations, additional simulations are performed for 35 ns in the NPT ensemble. These simulation trajectories are used to characterize dynamic heterogeneity by means of local segmental mobilities, as described in Sections 2.2.2 and 2.2.3. Finally, systems at 500 K and all compositions except $x^{\text{(PEO)}}=0.0$, and systems at 360 K and $x^{\text{(PEO)}}=0.6$ to $1.0$, are simulated for an additional 100 ns to enable the Rouse mode analysis described in Section 2.2.4. For $x^{\text{(PEO)}}=0.0$ at 500 K, the simulation is extended by 200 ns to improve statistical sampling.
We note that our system preparation protocol yields amorphous systems across all compositions and temperatures. As such, the following analysis pertains to well-mixed systems without the influence of crystallization, which is not accessible on simulation timescales. While semicrystallinity is an important factor for material properties, its absence here removes a potential confounding factor, allowing a clearer focus on the fundamental dynamical coupling between components.

### v2

@@ Abstract
The emergence of spatially variable local dynamics, or dynamic heterogeneity, is common in multicomponent polymer systems. Although often attributed to differences in the intrinsic dynamics of each component, the molecular origin of their coupling and its dependencies remain unclear. Here, we use molecular dynamics simulations of polyethylene oxide (PEO)/polymethyl methacrylate (PMMA) blends, across the full range of compositions and multiple thermal regimes, to characterize local fluctuations and sub-chain relaxations for both PEO and PMMA. By constructing probability distributions of local composition and computing entropic measures, we connect nanoscale heterogeneity to differences in mobility between PEO and PMMA, extending beyond mean‐field treatments. While PMMA segmental fluctuations in blends broadly align with $T_{\text{g}}$‐equivalent neat PMMA systems, PEO exhibits enhanced mobility correlated with increased free volume and broader, more diverse local compositions upon blending. Rouse‐mode analysis, used to probe relaxation dynamics over different length scales, shows that PEO relaxation approaches neat‐like behavior in PEO‐rich domains, whereas PMMA relaxation accelerates uniformly across all mode numbers. Given the local mobility enhancement of PMMA by PEO, this uniform shift suggests a nanoscale facilitation process that extends PEO’s influence beyond its immediate environment. These findings link the statistics of local compositional heterogeneity to dynamic asymmetry across length scales, provide physical insight into the behavior of this archetypal blend system, and establish a framework for analyzing dynamic coupling in others.
@@
Princeton University] Department of Chemical and Biological Engineering, Princeton University, Princeton, NJ, 08544 USA
@@ 1 Introduction
Dynamic heterogeneity is known to be significant in blends of polyethylene oxide (PEO) and polymethyl methacrylate (PMMA). While these polymers are miscible,22, 23, 24, 25 their glass transition temperatures ($T_{\text{g}}$) differ by approximately 180 K.26, 27, 28 This dissimilarity gives rise to pronounced differences in segmental dynamics when the polymers are blended, with relaxation times differing by up to 12 orders of magnitude.26, 29 Such contrast in mobilities has implications for applications like solid polymer electrolytes,30, 31, 32, 33, 34 where ionic conductivity is sensitive to local segmental fluctuations.35, 36, 37 More broadly, PEO/PMMA blends serve as model systems for investigating dynamic coupling between components at the nanoscale, with signatures of heterogeneity accessible in both simulations and experiments.29, 38, 39, 40
@@
Numerous experimental studies have examined the dynamics of PEO and PMMA in blends under varying conditions. Quasi-elastic neutron scattering (QENS) on blends containing up to 30 wt% PEO has shown that the segmental mobility of PMMA, on length scales up to 11 Å, is primarily governed by the temperature difference between the system and the glass transition temperature of the blend, (i.e., $T-T_{\text{g}}$).41, 42, 43 In contrast, PEO dynamics are strongly influenced by interactions with PMMA, exhibiting distinct behavior at short (<1 nm) and longer (>1 nm) length scales.44, 38 At short length scales, QENS,45, 39 nuclear magnetic resonance (NMR),26, 29 and neutron spin echo (NSE) spectroscopy46 reveal narrowing distributions of segmental relaxation times across a range of compositions and temperatures. These effects are attributed to self-concentration phenomena45 and local confinement by the stiffer PMMA matrix, with a dependence on local free volume.26, 29, 39
@@
On longer length scales, PEO dynamics have been investigated using infrared dichroism and birefringence,47 QENS,44, 39 and NSE.38 These measurements consistently indicate a pronounced slowdown of PEO segmental motion in blends with low PEO content.47, 44, 38, 39 Unlike the behavior observed at shorter scales, the dynamics at these larger scales are characterized by a broad distribution of relaxation times,39 which is largely attributed to long-range concentration fluctuations.47 Collectively, these observations highlight the presence of dynamic heterogeneity across multiple length scales in PEO/PMMA blends and suggest several underlying mechanisms; however, direct inference of the molecular-level phenomena remains elusive through experimental characterization alone.
@@
To complement experimental observations, several theoretical models have been developed to explain dynamic heterogeneity in PEO/PMMA blends. The Lodge-McLeish (LM) model, also referred to as the chain connectivity model, quantifies the influence of local self-concentration on polymer dynamics.48 The LM model predicts the segmental dynamics of each polymer component using the concentration within a cooperative volume centered on a monomer. The relevant length scale used to determine the size of the cooperative volume is the Kuhn length of each polymer component. It successfully predicts PEO relaxation times in blends containing 10-30 wt% PEO taken from QENS measurements for large spatial scales (qq=0.69 Å-1, approximately 18 Å) but breaks down for smaller spatial scales (qq=1.3 Å-1, approximately 10 Å).39 When the self-concentration is allowed to vary as a fitting parameter, the LM model reasonably fits PEO relaxation times measured by QENS39 and NMR,29 as well as terminal PMMA relaxation times.49 However, allowing the self-concentration to vary obfuscates the theoretical foundations of and insights from the model. Extensions of the LM model incorporate concentration fluctuations to quantitatively predict relaxation times in polymer blends.50, 51, 52, 53 These studies demonstrate that concentration fluctuations are necessary to accurately capture both the peak and width of relaxation time spectra, particularly at temperatures below the blend $T_{\text{g}}$. The correlation volume within which concentration fluctuations predict relaxation times is found to be on the order of the Kuhn length, supporting the assumptions of the LM model.
[…]
@@
Finally, a coupling model56, 57, 58, 59, 60 and the generalized Langevin equation framework61, 62, 63 have been developed to describe a crossover time that separates single-chain PEO dynamics at short length scales from many-chain coupled PEO dynamics at long length scales. This framework has been used to explain unexpected phenomena and properties that arise specifically in dynamically asymmetric miscible polymer blends, such as different $T_{\text{g}}$ values for each component and the breakdown of the time-temperature superposition.58 Additionally, the coupling model closely captures PEO segmental dynamics for experimental QENS measurements using a momentum transfer $q$ value between 1 and 2 Å-1 in blends with 10-30 wt% PEO.39 These theories qualitatively capture and rationalize trends in blended PEO dynamics up to mesoscale length scales. However, like with experimental observations, they do not conclusively illustrate the molecular-level phenomena that contribute to dynamic heterogeneity.
@@
Molecular dynamics (MD) simulations have helped to elucidate the microscopic origins of dynamic heterogeneity in polymer blends under certain conditions. Analysis of all-atom simulations with blend compositions of 10-30 wt% PEO suggest the presence of multiple populations of PEO dynamics in blends which may be caused by confinement effects of PEO in a rigid matrix of PMMA. This is deduced from van Hove self-correlation functions of hydrogen atoms from PEO, which show a double-peak structure at 400 K with an unmoving second peak44 and PEO relaxation times that are broadly distributed.64 Monitoring mean-square deviations of atoms in PEO and PMMA also suggests highly disparate local cage sizes, as the magnitude of distance traveled in the ballistic regime is much higher for PEO than PMMA.44, 64 Additionally, PEO segmental relaxation times have been found to be more stretched upon blending than those of PMMA.64 Rouse analyses based on simulations of 20 wt% PEO have illustrated the non-Gaussianity of the distribution of PEO atomic displacements in blends65, suggesting that local PEO segmental motion for wavelengths on the order of the size of a monomer is not strongly affected by blending but is more strongly affected at larger wavelengths.40 These results provide molecular-level insights into dynamic heterogeneity in PEO/PMMA blends, albeit using a varied set of compositions, temperatures, and characterization methods.
@@
In this work, we use atomistic MD simulations to systematically characterize how dynamic heterogeneity manifests at the nanoscale in PEO/PMMA blends. A central objective is to examine how the strength of dynamic coupling between the two components depends on temperature and blend composition. Although prior studies have explored the dynamic behavior of PEO/PMMA blends, inconsistencies in system specifications (e.g., blend composition, molecular weight, and force field) and thermodynamic conditions have made it difficult to identify clear trends. To address this, we comprehensively investigate behavior over the full range of blend compositions at temperatures above, between, and below the simulated apparent apparent $T_{\text{g}}$ of both polymers. Dynamics are characterized by means of both local segmental fluctuations as well as segmental relaxation timescales. We find that blending induces asymmetric and composition-dependent changes in both local and collective dynamics, governed by temperature and local composition. In particular, the free volume and the diversity in local composition surrounding polymer units are found to correlate strongly with differences in segmental mobility between PEO and PMMA. These findings help to clarify what PEO-PMMA interactions influence dynamic heterogeneity and motivate further investigation into their generality and implications for macroscopic properties and functional performance.
@@ 2.1.1 General simulation protocols
All MD simulations are performed using LAMMPS (ver 29, Sep 2021).66 Systems are modeled using the all-atom optimized potentials for liquid simulations (OPLS-AA)67 force field. Real-space non-bonded interactions are truncated at 12 Å. Long-range electrostatics are handled using the particle-particle-particle-mesh Ewald summation method with a $10^{-4}$ convergence accuracy.68, 69 Equations of motion are integrated using a velocity-Verlet integration scheme and a 1 fs timestep. Periodic boundary conditions are used in all three dimensions. Temperature and pressure are controlled using a Nosé-Hoover thermostat and barostat with damping constants of 100 fs and 2000 fs, respectively.
@@ 2.1.2 Polymer chain generation
Each PEO chain consists of 75 monomers, and each PMMA chain consists of 33 monomers. Here and throughout the text, the term monomer is used to refer to a single constitutional repeat unit of the polymer. As a result, chains for both polymers possess molecular weights of approximately 3,300 g/mol. Syndiotactic PMMA was used in all simulations. This choice avoids the need to stochastically generate atactic sequences. Moreover, syndiotactic and atactic PMMA have been previously noted to exhibit similar $T_{\text{g}}$,70 which may suggest similar dynamical characteristics.
Initial chain configurations are generated to approximate the expected relationship for the mean-squared end-to-end distance in that $\langle R^{2}\rangle=L\ell_{k}$, where $\ell_{k}$ is the Kuhn length of the polymer and $L$ is the statistical contour length. In practice, rather than $L$ we use $L_{\text{ext}}\approx\sum_{i}l_{i}$, where $L_{\text{ext}}$ is the length of a fully extended polymer chain approximated by the summation of individual bond lengths $l_{i}$. The individual $l_{i}$ are obtained from repeat units, following geometry optimization in Avogadro 1.2.0 with the steepest descent algorithm and the UFF force field;71, 72 this yields $L$ values of 326.1 Å for PEO and 102.5 Å for PMMA. While this approach will initially yield somewhat more extended conformations than targeted as $L_{\text{ext}}>L$, this approximation is used only for initialization, and subsequent equilibration mitigates initial bias in chain dimensions. The Kuhn lengths used are 8.2 Å for PEO and 13.8 Å for PMMA.73 Chains are constructed by sequentially adding monomers, with each new monomer positioned based on a randomly sampled dihedral angle. For PEO, dihedral angles are drawn from a uniform distribution between 0.65-0.65 and $0.65$ radians; for PMMA, the range is 0.4-0.4 to $0.4$ radians. For each system described below, this process is repeated until the required number of independent chain configurations within a threshold of 5 Å of the target $\langle R^{2}\rangle$ is obtained. These configurations are used as described in Section 2.1.3. The contour length of each polymer is estimated by using geometry optimization in Avogadro 1.2.0 to measure the distance between two monomers on a fully extended chain and multiplying by the total number of monomers. This yields estimated contour lengths of 273 Å for PEO and 92 Å for PMMA. Dividing these values by the respective Kuhn lengths gives 33.3 and 6.7 Kuhn steps for PEO and PMMA chains, respectively.
@@ 2.1.3 System preparation
For each system, an initial configuration is generated by randomly placing each pre-constructed chain into a 60 Å $\times$ 60 Å $\times$ 60 Å simulation box with random positions and orientations. A brief energy minimization is then performed where the system is simulated in the microcanonical (NVE) ensemble with a constrained maximum distance of 0.005 Å moved per timestep; this minimization takes place for 0.05 ns. After, the constrained maximum distance is increased to 0.1 Å  and the system is simulated for 0.5 ns. This procedure resolves unfavorable atomic overlaps introduced by the random packing procedure. Following minimization, initial particle velocities are randomly generated from a uniform distribution over the interval [0.5,0.5][-0.5,0.5], rescaled such that temperature as estimated from the kinetic energy corresponds to 300 K, and finally shifted to remove any net linear momentum in the simulation cell. The system then undergoes 1 ns of simulation in the NVE ensemble, followed by 1 ns in the canonical (NVT) ensemble at 300 K. Subsequently, a barostat is introduced to maintain the pressure at 1 atm, and the system is heated from 300 K to 700 K at a rate of 80 K/ns. After reaching 700 K, the system is equilibrated in the isothermal–isobaric (NPT) ensemble for 50 ns. Finally, the system is cooled from 700 K to 100 K at a rate of 10 K/ns at constant pressure.
The cooling trajectories are used for analysis of glass transition temperatures (see Section 2.2.1). Simulation configurations are also recorded specifically around the temperatures of 500 K, 360 K, and 220 K. From these configurations, additional simulations are performed for 35 ns in the NPT ensemble. These simulation trajectories are used to characterize dynamic heterogeneity by means of local segmental mobilities, as described in Sections 2.2.2 and 2.2.3. Finally, additional simulation is performed for systems at 500 K to enable the Rouse mode analysis described in Section 2.2.5; for $x^{\text{(PEO)}}=1.0$, trajectories are extended by 100 ns, and for all other systems, trajectories are extended by 200 ns, with the difference being due to slower relaxation.
All blends are prepared well-mixed and remain so for the duration of the simulation. All systems, including neat PEO and PMMA, remain amorphous throughout our simulations. Crystallization is neither observed nor expected on the timescale of our simulations. Partially by consequence, demixing is also not observed across the full composition range. As a result, the data presented are representative of hypothetical miscible systems. The simulations do not establish miscibility for PEO/PMMA blends in this regime but rather explore the consequences of it in a simulation setting.